Jodi poured herself a cold soda that had an initial temperature of 41
degreesF and immediately went outside to sunbathe where the temperature was a steady 100degreesF. After 8 minutes the temperature of the soda was 44
degreesF. Jodi had to run back into the house to answer the phone. What is the expected temperature of the soda after an additional 5 ​minutes?
(Do not round until the final answer. Then round to the nearest tenth as​ needed.)

Answer by stanbon(75887)

Jodi poured herself a cold soda that had an initial temperature of 41
degreesF and immediately went outside to sunbathe where the temperature was a steady 100degreesF. After 8 minutes the temperature of the soda was 44
degreesF. Jodi had to run back into the house to answer the phone. What is the expected temperature of the soda after an additional 5 ​minutes?
(Do not round until the final answer. Then round to the nearest tenth as​ needed.)
-----
soda gained 3 degrees in 8 minutes or 3/8 degrees per minute
-----
In 5 minutes it would gain 5(3/8) = 15/8 degrees
Ans: 44 degrees + (15/8)degrees = 45.875 degrees
--------------
Cheers,
Stan H.
----------